# LYT-NET: Lightweight YUV Transformer-based Network for Low-light Image Enhancement

### v6

@@ I Introduction
Earlier LLIE methods [1] primarily relied on frequency decomposition [2, 3, 4], histogram equalization [5, 6, 7], and Retinex theory [8, 9, 10, 11, 12]. With the rapid advancement of deep learning, various CNN architectures [13, 14, 15, 16, 17, 18, 19, 20, 21, 22] have been shown to outperform traditional LLIE techniques. Based on Retinex theory, Retinex-Net [13] integrates Retinex decomposition with an original CNN architecture, while Diff-Retinex [16] proposes a generative framework to further address content loss and color deviation caused by low light.
Figure 1: Our model (LYT-Net) delivers SOTA performance in LLIE task, while maintaining the best computational efficiency (results are plotted on LOL dataset [13]).
@@
More recently, Vision Transformers (ViTs) [25] have demonstrated significant effectiveness in various computer vision tasks [26, 27, 28, 29, 30], largely due to the self-attention (SA) mechanism. Despite these advancements, the application of ViTs to low-level vision tasks remains relatively underexplored. Only a few LLIE (Low-Light Image Enhancement)-ViT-based strategies have been introduced in the recent literature [31, 32, 33]. Uformer [31] is based on the classical UNet architecture, where the convolution layers are replaced with Transformer blocks while maintaining the hierarchical encoder-decoder structure and skip connections. Restormer [33], on the other hand, introduces a multi-Dconv head transposed attention (MDTA) block, replacing the vanilla multi-head self-attention.
@@ II-A Channel-wise Denoiser Block
The CWD Block employs a U-shaped network with MHSA as the bottleneck, integrating convolutional and attention-based mechanisms. It includes multiple conv3$\times$3 layers with varying strides and skip connections, facilitating detailed feature capture and denoising.
It consists of a series of four conv3$\times$3 layers. The first conv3$\times$3 has strides of $1$ for feature extraction. The other three conv3$\times$3 layers have strides of $2$, helping with capturing features at different scales. The integration of the attention bottleneck enables the model to capture long-range dependencies, followed by upsampling layers and skip connections to reconstruct and facilitate the recovery of spatial resolution.
@@ II-B Multi-headed Self-attention Block
In our updated simplified transformer architecture, the input feature $\mathbf{F}_{\text{in}}\in\mathbb{R}^{H\times W\times C}$ is first linearly projected into query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$) components through bias-free fully connected layers. The linear projections maintain the original input dimensionality.
$\small\mathbf{Q}=\mathbf{X}\mathbf{W}_{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}_{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}_{\mathbf{V}},~{}\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{HW\times C}\vspace{0.5mm}$ (1)
Next, these projected features are split into $k$ heads:
$\small\mathbf{X}=[\mathbf{X}_{1},\mathbf{X}_{2},\cdots,\mathbf{X}_{k}],~{}\mathbf{X}_{i}\in\mathbb{R}^{HW\times d_{k}},~{}d_{k}=\frac{C}{k},i=\overline{1,k}\vspace{-0.5mm}$ (2)
@@
$\small\text{Attention}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\text{T}}}{\sqrt{d_{k}}}\right)\!\times\!\mathbf{V}_{i}\vspace{-0.5mm}$ (3)
@@ II-C Multi-stage Squeeze & Excite Fusion Block
These operations compress the feature map into a reduced descriptor (the squeezing operation) to capture essential details, and then re-expand it (the excitation operation) to restore the full dimensions while emphasizing the most relevant features.
$\small\mathbf{S}_{\text{reduced}}=\text{ReLU}(\mathbf{W}_{1}\cdot\text{GlobalPool}(\text{LayerNorm}(\mathbf{F}_{\text{in}})))\vspace{-0.5mm}$ (4)
@@
$\small\mathbf{F}_{\text{out}}=\text{DWConv}(\text{LayerNorm}(\mathbf{F}_{\text{in}}))\cdot\mathbf{S}_{\text{expanded}}+\mathbf{F}_{\text{in}}\vspace{-0.5mm}$ (6)
  Complexity LOL-v1 LOL-v2-real LOL-v2-syn
Methods FLOPS (G) Params (M) PSNR SSIM PSNR SSIM PSNR SSIM
  SID [34] 13.73 7.76 14.35 0.436 13.24 0.442 15.04 0.610
3DLUT [35] 0.075 0.59 21.35 0.585 20.19 0.745 22.17 0.854
DeepUPE [15] 21.10 1.02 14.38 0.446 13.27 0.452 15.08 0.623
DeepLPF [14] 5.86 1.77 15.28 0.473 14.10 0.480 16.02 0.587
UFormer [31] 12.00 5.29 16.36 0.771 18.82 0.771 19.66 0.871
RetinexNet [13] 587.47 0.84 18.92 0.427 18.32 0.447 19.09 0.774
Sparse [36] 53.26 2.33 17.20 0.640 20.06 0.816 22.05 0.905
EnGAN [24] 61.01 114.35 20.00 0.691 18.23 0.617 16.57 0.734
FIDE [37] 28.51 8.62 18.27 0.665 16.85 0.678 15.20 0.612
KinD [17] 34.99 8.02 20.86 0.790 14.74 0.641 13.29 0.578
Restormer [33] 144.25 26.13 26.68 0.853 26.12 0.853 25.43 0.859
MIRNet [20] 785 31.76 26.52 0.856 27.17 0.865 25.96 0.898
SNR-Net [22] 26.35 4.01 26.72 0.851 27.21 0.871 27.79 0.941
Retinexformer [32] 15.57 1.61 27.14 0.850 27.69 0.856 28.99 0.939
LYT-Net 3.49 0.045 27.23 0.853 27.80 0.873 29.38 0.940
TABLE I: Quantitative results on LOL datasets. Best results are in red, second best are in blue.
@@ II-D Loss Function
$\small\mathbf{L}=\mathbf{L}_{\text{S}}+\alpha_{1}\mathbf{L}_{\text{Perc}}+\alpha_{2}\mathbf{L}_{\text{Hist}}+\alpha_{3}\mathbf{L}_{\text{PSNR}}+\alpha_{4}\mathbf{L}_{\text{Color}}+\alpha_{5}\mathbf{L}_{\text{MS-SSIM}}\vspace{-0.5mm}$ (7)
@@ III Results and Discussion
Implementation details: The implementation of LYT-Net utilizes the TensorFlow framework. The ADAM Optimizer ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) is employed for training over $1000$ epochs. The initial learning rate is set to $2\times 10^{-4}$ and gradually decays to $1\times 10^{-6}$ following a cosine annealing schedule, aiding in optimization convergence and avoiding local minima. The hyperparameters of the hybrid loss function are set as: $\alpha_{1}\!\!=\!\!0.06$, $\alpha_{2}\!\!=\!\!0.05$, $\alpha_{3}\!\!=\!\!0.5$, $\alpha_{4}\!\!=\!\!0.0083$, and $\alpha_{5}\!\!=\!\!0.25$.
LYT-Net is trained and evaluated on three versions of the LOL dataset: LOL-v1, LOL-v2-real, and LOL-v2-synthetic. The corresponding training/testing splits are $485:15$ for LOL-v1, $689:100$ for LOL-v2-real, and $900:100$ for LOL-v2-synthetic.
@@
Quantitative results: The proposed method is compared to SOTA LLIE techniques, as shown in Table I, focusing on two aspects: quantitative performance on the LOL datasets (LOL-v1, LOL-v2-real, LOL-v2-synthetic) and model complexity.
As shown in Table I, LYT-Net consistently outperforms the current SOTA methods across all versions of the LOL dataset in terms of both PSNR and SSIM. Additionally, LYT-Net is highly efficient, requiring only 3.49G FLOPS and utilizing just 0.045M parameters, which gives it a significant advantage over other SOTA methods that are generally much more complex. The only exception is 3DLUT[35], which is comparable to our approach in terms of complexity. However, LYT-Net clearly surpasses the 3DLUT method in both PSNR and SSIM. This combination of strong performance and low complexity highlights the overall effectiveness of LYT-Net.
@@
Previous methods, such as KiND[17] and Restormer[33], exhibit color distortion issues, as shown in Fig. 3. Additionally, several algorithms (e.g. MIRNet[20], and SNR-Net[22]) tend to produce over- or under-exposed areas, compromising image contrast while enhancing luminance. Similarly, Fig. 4 demonstrates that SRIE [39], DeHz [40], and NPE [41] result in a loss of contrast. In general, our LYT-Net is highly effective at improving visibility and enhancing low-contrast or poorly lit areas, while efficiently eliminating noise without introducing spots or artifacts.
@@ IV Ablation Study
  Y-CWD UV-CWD MSEF Params PSNR
40238 26.62
44377 26.99
48516 26.76
40784 26.78
44923 27.23
49062 27.02
@@
The ablation study is conducted on the LOLv1 dataset, using PSNR as the quantitative metric, and evaluates the impact of the CWD and MSEF blocks. In the YUV decomposition, applying CWD to the Y-channel (used as the illumination map) results in the retention of lighting artifacts, leading to performance degradation compared to pooling operations and interpolation-based upsampling, which smooth the illumination for better outcomes. However, CWD enhances the chrominance channels (U and V), preserving detail without introducing noise. Moreover, the MSEF block consistently boosts performance across all CWD combinations, improving PSNR by 0.16, 0.24, and 0.26 dB, respectively, while increasing the parameter count by only 546.
@@ V Conclusions
We introduce LYT-Net, an innovative lightweight transformer-based model for enhancing low-light images. Our approach utilizes a dual-path framework, processing chrominance and luminance separately to improve the model’s ability to manage illumination adjustments and restore corrupted regions. LYT-Net integrates multiple layers and modular blocks, including two unique components—Channel-Wise Denoiser (CWD) and Multi-Stage Squeeze & Excite Fusion (MSEF)—as well as the traditional Vision Transformer (ViT) block with Multi-Headed Self-Attention (MHSA). A comprehensive qualitative and quantitative analysis demonstrates that LYT-Net consistently outperforms SOTA methods on all versions of the LOL dataset in terms of PSNR and SSIM, while maintaining high computational efficiency.

### v7

@@ I Introduction
Earlier LLIE methods [1] primarily relied on frequency decomposition [2, 3], histogram equalization [4, 5], and Retinex theory [6, 7, 8, 9]. With the rapid advancement of deep learning, various CNN architectures [10, 11, 12, 13, 14, 15, 16, 17, 18, 19] have been shown to outperform traditional LLIE techniques. Based on Retinex theory, Retinex-Net [10] integrates Retinex decomposition with an original CNN architecture, while Diff-Retinex [12] proposes a generative framework to further address content loss and color deviation caused by low light.
Figure 1: Our model delivers SOTA performance in LLIE task, while maintaining computational efficiency (results are plotted on LOL dataset [10]).
@@
More recently, Vision Transformers (ViTs) [22] have demonstrated significant effectiveness in various computer vision tasks [23, 24, 25, 26], largely due to the self-attention (SA) mechanism. Despite these advancements, the application of ViTs to low-level vision tasks remains relatively underexplored. Only a few LLIE-ViT-based strategies have been introduced in the recent literature [27, 28, 29, 30]. Restormer [29], on the other hand, introduces a multi-Dconv head transposed attention (MDTA) block, replacing the vanilla multi-head self-attention.
@@
Diffusion models have emerged as a powerful approach for LLIE, leveraging their ability to learn complex data distributions through a simulated forward process [31, 32, 33].
@@ II-A Channel-wise Denoiser Block
[…]
@@ II-B Multi-headed Self-attention Block
In our updated simplified transformer architecture, the input feature $\mathbf{F}_{\text{in}}\in\mathbb{R}^{H\times W\times C}$ is first linearly projected into query ($\mathbf{Q}$), key ($\mathbf{K}$), and value ($\mathbf{V}$) components through bias-free fully connected layers. The linear layers use parameter $D$ to determine projection head dimensionality.
$\small\mathbf{Q}=\mathbf{X}\mathbf{W}_{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}_{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}_{\mathbf{V}},~\mathbf{Q},\mathbf{K},\mathbf{V}\in\mathbb{R}^{HW\times D}$ (1)
where $\mathbf{W}_{\mathbf{Q}},\mathbf{W}_{\mathbf{K}},\mathbf{W}_{\mathbf{V}}$ are fully connected layer weights. Next, these projected features are split into $k$ heads as such:
$\small\mathbf{X}=[\mathbf{X}_{1},\mathbf{X}_{2},\cdots,\mathbf{X}_{k}],~\mathbf{X}_{i}\in\mathbb{R}^{HW\times d_{k}},~d_{k}=\frac{D}{k},i=\overline{1,k}\vskip-1.42262pt$ (2)
@@
$\small\text{Attention}(\mathbf{Q}_{i},\mathbf{K}_{i},\mathbf{V}_{i})=\text{Softmax}\left(\frac{\mathbf{Q}_{i}\mathbf{K}_{i}^{\text{T}}}{\sqrt{d_{k}}}\right)\!\times\!\mathbf{V}_{i}\vskip-1.42262pt$ (3)
@@ II-C Multi-stage Squeeze & Excite Fusion Block
These operations compress the feature map into a reduced descriptor (the squeezing operation) to capture essential details, and then re-expand it (the excitation operation) to restore the full dimensions, emphasizing the most relevant features.
$\small\mathbf{S}_{\text{reduced}}=\text{ReLU}(\mathbf{W}_{1}\cdot\text{GlobalPool}(\text{LayerNorm}(\mathbf{F}_{\text{in}})))$ (4)
@@
$\small\mathbf{F}_{\text{out}}=\text{DWConv}(\text{LayerNorm}(\mathbf{F}_{\text{in}}))\cdot\mathbf{S}_{\text{expanded}}+\mathbf{F}_{\text{in}}$ (6)
Consequently, the MSEF block acts as a multilayer perceptron capable of performing efficient feature extraction on self-attended and denoised chrominance features, enabling high-quality restoration with minor parameter count increase.
                 Complexity LOL-v1 LOL-v2-real LOL-v2-syn       SDSD
Methods FLOPS (G) Params (M) PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
  3DLUT              [34] 0.075 0.59 21.35 0.585 20.19 0.745 22.17 0.854 21.78 0.652
UFormer [27] 12.00 5.29 16.36 0.771 18.82 0.771 19.66 0.871 23.51 0.804
RetinexNet [10] 587.47 0.84 18.92 0.427 18.32 0.447 19.09 0.774 20.90 0.623
Sparse [35] 53.26 2.33 17.20 0.640 20.06 0.816 22.05 0.905 24.27 0.834
EnGAN [21] 61.01 114.35 20.00 0.691 18.23 0.617 16.57 0.734 20.06 0.610
FIDE [36] 28.51 8.62 18.27 0.665 16.85 0.678 15.20 0.612 22.31 0.644
LYT-Net (Ours) 3.49 0.045 22.38 0.826 21.83 0.849 23.78 0.921 28.42 0.877
                    KinD      [13] 34.99 8.02 20.86 0.790 14.74 0.641 13.29 0.578 21.96 0.663
Restormer [29] 144.25 26.13 26.68 0.853 26.12 0.853 25.43 0.859 25.23 0.815
DepthLux [37] - 9.75 26.06 0.793 26.16 0.794 28.69 0.920 - -
ExpoMamba [38] - 41 25.77 0.860 28.04 0.885 - - - -
MIRNet [16] 785 31.76 26.52 0.856 27.17 0.865 25.96 0.898 25.76 0.851
SNR-Net [18] 26.35 4.01 26.72 0.851 27.21 0.871 27.79 0.941 29.05 0.880
KAN-$\mathcal{T}$ [39] - 2.80 26.66 0.854 28.45 0.884 28.77 0.939 - -
Retinexformer [28] 15.57 1.61 27.14 0.850 27.69 0.856 28.99 0.939 29.81 0.887
LYT-Net (Ours) 3.49 0.045 27.23 0.853 27.80 0.873 29.38 0.940 28.42 0.877
TABLE I: Quantitative results on LOL datasets. Best results are in red, second best are in blue. Highlighted cells show results with GT-Mean gamma correction [13], which is widely used on the LOL datasets..
@@ II-D Loss Function
$\small\mathbf{L}=\mathbf{L}_{\text{S}}+\alpha_{1}\mathbf{L}_{\text{Perc}}+\alpha_{2}\mathbf{L}_{\text{Hist}}+\alpha_{3}\mathbf{L}_{\text{PSNR}}+\alpha_{4}\mathbf{L}_{\text{Color}}+\alpha_{5}\mathbf{L}_{\text{MS-SSIM}}\vskip-1.42262pt$ (7)
@@ III Results and Discussion
Implementation details: The implementation of LYT-Net utilizes the TensorFlow framework. The ADAM Optimizer ($\beta_{1}=0.9$ and $\beta_{2}=0.999$) is employed for training over $1000$ epochs. The initial learning rate is set to $2\times 10^{-4}$ and gradually decays to $1\times 10^{-6}$ following a cosine annealing schedule, aiding in optimization convergence and avoiding local minima. The hyperparameters of the hybrid loss function are set as: $\alpha_{1}\!\!=\!\!0.06$, $\alpha_{2}\!\!=\!\!0.05$, $\alpha_{3}\!\!=\!\!0.0083$, $\alpha_{4}\!\!=\!\!0.25$, and $\alpha_{5}\!\!=\!\!0.5$.
LYT-Net is trained and evaluated on: LOL-v1, LOL-v2-real, and LOL-v2-synthetic. The corresponding training/testing splits are $485:15$ for LOL-v1, $689:100$ for LOL-v2-real, and $900:100$ for LOL-v2-synthetic.
@@
Quantitative results: The proposed method is compared to SOTA LLIE techniques, as shown in Table I, focusing on several aspects: quantitative performance on the LOL (LOL-v1, LOL-v2-real, LOL-v2-synthetic) and SDSD [40] datasets, and model complexity.
As shown in Table I, LYT-Net consistently outperforms the current SOTA methods across all versions of the LOL dataset in terms of both PSNR and SSIM. Additionally, LYT-Net is highly efficient, requiring only 3.49G FLOPS and utilizing just 0.045M parameters, which gives it a significant advantage over other SOTA methods that are generally much more complex. The only exception is 3DLUT[34], which is comparable to our approach in terms of complexity. However, LYT-Net clearly surpasses the 3DLUT method in both PSNR and SSIM. This combination of strong performance and low complexity highlights the overall effectiveness of LYT-Net. On SDSD, where images are high resolution, our method shows limitations due to its significantly low parameter count. However, by utilizing a deeper variant of LYT-Net, we expect that performance increases accordingly.
@@
Previous methods, such as KiND[13] and Restormer[29], exhibit color distortion issues, as shown in Fig. 3. Additionally, several algorithms (e.g. MIRNet[16], and SNR-Net[18]) tend to produce over- or under-exposed areas, compromising image contrast while enhancing luminance. Similarly, Fig. 4 demonstrates that SRIE [42], DeHz [43], and NPE [44] result in a loss of contrast. In general, our LYT-Net is highly effective at enhancing visibility and low-contrast or poorly lit areas, while efficiently eliminating noise without introducing artifacts.
@@ IV Ablation Study
The ablation study is conducted on the LOLv1 dataset, using PSNR and CIEDE2000 [45] as quantitative metrics, and evaluates the impact of the CWD and MSEF blocks. In the YUV decomposition, applying CWD to the Y-channel (used as the illumination map) results in the retention of lighting artifacts, leading to performance degradation compared to pooling operations and interpolation-based upsampling, which smoothen the illumination for better and more uniform lighting. However, CWD enhances the chrominance channels (U and V), preserving detail without introducing noise. Moreover, the MSEF block consistently boosts performance across all CWD combinations, improving PSNR by 0.16, 0.24, and 0.26 dB, respectively, only increasing the parameter count by 546.
[…]
40238 26.62 6.3087
44377 26.99 6.0148
48516 26.76 6.1975
40784 26.78 6.1816
44923 27.23 5.8242
49062 27.02 5.9910
@@ V Conclusions
We introduce LYT-Net, an innovative lightweight transformer-based model for enhancing low-light images. Our approach utilizes a dual-path framework, processing chrominance and luminance separately to improve the model’s ability to manage illumination adjustments and restore corrupted regions. LYT-Net integrates multiple layers and modular blocks, including two unique CWD and MSEF — as well as the traditional ViT block with MHSA. A comprehensive qualitative and quantitative analysis demonstrates that LYT-Net consistently outperforms SOTA methods on all versions of the LOL dataset in terms of PSNR and SSIM, while maintaining high computational efficiency.
Acknowledgement: Part of this research is supported by the ”Romanian Hub for Artificial Intelligence – HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906.